(25%) Problem 4: A uniform rod of mass M and length L is free to swing back and forth by pivoting a distance x from its centea,b,c please
Answer #1

Here we apply concept of moment of inertia of solid body and parallel axis theorem as well as concept of simple harmonic motion of solid body.

Por May = Length = 2, M=2.4kg L=1-6m Position of pivot from center, x = 0.43m de Moment of Inertia of t e e red about centerTime Period, T = 24 - 25 le To max IT=an M (22/2+222 =21/(22 +1282) J Mga v 12g & s T=21 [cl•6m)?+1266.73m)2] 12 (9.8m/s2)60:+ना +12x4 (1292) 27 dT = 4+2_d (12+12x2) ____dx (129X) 20 + 12 ---+12 =0